SYMMETRIC RELAY BRIDGE SYSTEM

1. Openings

1♣: 16+ points any or distributional equivalent. 1♢: 10 - 15 points; either 5-4 primary minor, or 4-4-4-1 distribution in any suit, or 5-5 in the minors. 1♡/1♠: 10 - 15 points; 5-card suit. 1 NT : 12 - 15 points; balanced hand. 2♣/2♢: 10 - 15 points; six-card suit. 2♡/2♠: 0 - 9 points; 5^{+}-card suit. 2 NT : 0 - 10 points; 5-5 in the minors. 3^{+} of a suit: natural and pretty weak. 3 NT: a good 4-major opener

2. Some responses

The responces are pretty natural except for 1♣ responces and the following ones: Responses to 1♢: They are pretty natural except for: 1♢ − 1♡ − 2♢: good support in hearts; 1♢ − 1♡ − 2♡: minimal support in hearts; 1♢ − 1♠ − 2♡: good support in spades; 1♢ − 1♠ − 2♠: minimal support in spades; 1♢ − 1♡/1♠ − 1 NT : three-card support in hearts/spades; 1♢ − 1♡ − 2♣: 5-4 or 5-5 in the minors; 1♢ − 1♠ − 2♣ : ♣&♢ or ♣&♡; 1♢ − 1♠ − 2♢ : ♢&♡. Responses to 1 NT : Again standard except for: 1 NT −2♠: which is either invitational or strong; 1 NT −2 NT: transfer to clubs; 1 NT −3♣ : transfer to diamonds. Responses to 2♣ : Raises are weak.

2♣ − 2♢ : balanced hand; invitational+. 2♣ − 2♡/2♠ : five-card suit; invitational+. 2♣ − 2 NT : a good invitational raise.

Responses to 2♢ :

Raises are weak. 2♢ − 2♡/2♠ : five-card suit; invitational+. 2♢ − 2 NT : balanced; the opener bids 3♣ if he has maximum hand, 3♢ for minimal. 2♢ − 3♣ : a good invitational raise.

3. Strong club relays

1♣ − 1♢: less than 8 points. The responses from the opener are either natural (jumps show at least 20 points) or artificial :

- 1♡ : 20^{+} any except for 20 − 23 with ♢;

- 1♠ : ♠& Another / ♣;

- 1 NT: 16 − 19 Semi-balanced;

- 2♣ : 16 − 23, ♢;

- 2♢ : 16 − 19, ♡;

- 2♡ : 5^{+}♡, 4♠;

- 2♠ : 16 − 19, ♠;

After 1♣−1♢−(1♡/1♠)−1 NT, this goes straight into relays as after 1♣. After the invite opener rebids 3♡ with minimum when responder can correct to 3♠/ pass. With maximum bid naturally.

This structure is on if opener has the opportunity to Double / bid a 1♡ opening else natural bidding with T/O doubles.

From here on the responder has at least 8 points.

1♣ − 1♡: 4+ spades, unbalanced.

- Relay of 1♠ -

- 1 NT: either a single suited hand in ♠ or a two suited hand in ♠&♣. It is followed by a relay of 2♣ and the answer is: 2♢ : ♠ or 2♡ : ♠&♣.

- 2♣: ♠&♢ - two suited hand.

- 2♢: three suited hand including spades. For further description see the end of the section.

- 2♡+ : ♠&♡ - two suited hand.

1♣ − 1♠ : either balanced (no 5 card suit or shortage), or ♣&♢ or ♣.

- Relay of 1 NT -

- 2♣ : ♣&♢

- 2♢ : ♣

- 2♡+: balanced. More precisely:

- 2♡ : 4♠ − 2♠ (R) - 2 NT : 44♡&♠, 3♣ : 4♣, 3♢ : 3433, 3♡ : 2443, 3♠/NT : 3442

- 2♠ : 4♡, < 4♠ - 2 NT (R) - 3♣ : 4♣, 3♢ : 3433, 3♡ : 2443, 3♠/NT : 3442

- 2 NT: 44♣&♢ − 3♣(R) − 3♢ : 2344, 3♡ : 3244

- 3♣ : 3334

- 3♢ : 3343

1♣ − 1 NT: ♡/♡&♣

- Relay of 2♣ -

SYMMETRIC RELAY BRIDGE SYSTEM 3

- 2♢: ♡ one suited hand.

- 2♡ : ♡&♣ two suited hand. 1♣ − 2♣ : ♡&♢ 1♣ − 2♢ : ♢ 1♣ − 2♡ : 3 suiter, short spades.

- Relay of 2♠ -

- 2 NT: 0445

- 3♣ : 0454

- 3♢ : 0544

- 3♡ : 1444 More about the sequence: 1♣ − 1♡ − 1♠(R) − 2♢ − 2♡(R):

- 2♠: high shortage (in hearts) / 4144

- 2 NT : middle shortage (in diamonds)

- 3♣ : 4414/4441

- 3♢ : 4450

- 3♡ : 4540

- 3♠/NT : 5440

4. The Schemes

4.1. One Suiters. After you’ve reached 2♢ − 2♡(R):

−2♠ − 2NT (R) − 3♣ = 6322/7222 (numeric split), 3♢^{+} = high shortage (as below) −2NT − 3♣(R) − 3♢+ as below with middle shortage −3♣ : 6232/6223 − R− numeric split. Low shortage goes direct −3♢ : 5332 −3♡ : 6331 −3♠/NT : 7(xx)1 −4♣ : 7330

4.2. Two Suiters. The scheme here is to first show your suits, then which is longer (or 5^{+}, 5^{+}) then pattern out much as for one suiters.

Having shown your suits and reached 2♣ − 2♢(R): −2♡: lower suit is longer (not 55^{+}) −2♠(R)− then as below −2♠ : 55^{+} From here on if bid directly, the higher suit is longer.

−2 NT: high shortage

−3♣: balanced (5422/7411) - R - numeric

−3♢ : 5431

−3♡ : 6421

−3♠/NT : 6430

−4♣ : 7420 (numeric order)

After 2♠ − R− −3♣ : high shortage −3♢ : 5611/6511/66(01) −3♡ : 5521 −3♠/NT : 5530 −4♣ : 5620 −4♢ : 6520

4.3. Relay Termination. The 3♠/NT things always happens in strong club openings.

3♠ / NT

This is triggered (A) by the relays stopping at 3♠, or (B) a bid of 3♡ if the relays have stopped lower.

With 8 − 11/15^{+} points you bid 3♠, with 12 − 14 you bid 3 NT. After either of these bids 4♢ is the Puppet Terminator (see below), 4♣/♡/♠/NT are RKB in partners longest/next longest/etc. suit (ties broken in order ♣/♢/♡/♠), see below for continuations. 3NT over 3♠ is to play.

With 15^{+} you bid 4♣^{+} over a 3NT sign off/ 4♠^{+} over a 4♢ puppet showing 5−, 6, 7 … controls after which new suits are RKB in length order again, with 5NT as an attempt to sign off in a small slam (puppets 6♣).

The 3♠ puppet

The major alternative to the scheme above is to go straight into RKB / sign off / make a quantative slam try. This isn’t always available (relays reach 3♠/NT) but normally is.

At this point 3♠ puppets 3NT after which :

4♣ is RKB for partner’s “third” longest suit,

4♢ is the puppet terminator (which triggers 4♡ bid),

4♡/4♠/4NT/5♣/5♢ are mild slam tries to be accepted by a decent 12 − 14 point hands. These set trumps and after 4♡ − NT keycards (controls 4−, 5, …) are shown when acccepting. After acceptance it is possible to play in a lower ranked suit slam (to facilitate 4−4 minor suit slams where you otherwise wish to play 4 of a major).

At the same time 3NT/4♣/4♢ direct are RKB for partner’s longest / second longest / shortest suits. If 3♡ is available then it is used for range as specified above. If 3♢ is available it acts as RKB for the longest suit with 4♣ via 3♠ keeping its place as RKB for the third longest suit and 4♣ direct being RKB for the shortest suit.

N.B. 3♢ here will nearly always be range RKB. Plenty of space to sort it all out though :)

If you aren’t giving range and stop early then 3♠ puppet is still on but all other are RKB in normal length order including 3♡.

Slightly more relay stuff

After a 3♠ bid showing 9 − 11/15^{+} when opener jumps straight into RKB and you have 15^{+} then with 2 keycards you jump straight to step 5^{+} spiral scanning.

With more you’re fine of course but with 0 or 1 it is not quite sure (0/1 is rather unlikely, esp. if opener is after keycards).

Also after a bid to show 15^{+} controls then if responder can bid 4♠ then that puppets 4 NT for sign offs. If the biddings reached 4♠ then use 4 NT as a last sign off attempt. If responder has bid 4 NT then there’s so many controls signing off would be silly. All others are RKB in length order.

If the shape showing has reached 4♢ then you can only play in game in one of responder’s 2 suits. Other bids are RKB in length order.

Really early relay breaks

A bid which isn’t the relay bid early one is a relay break. This is mainly done if you have a void and know the trump suit. As long as the break is 3♢^{+} it is handled as above. Step 1 and above agree partners last shown suit and show voids(in order ♣/♢/♡/♠ ). These dump you into cue bidding with serious 3 NT in majors. (Clearly only ever 3 voids to show :))

RKB and continuations

When the hand being relayed out has shown range via the 3♠/NT system then the system uses standard responses to RKB step 1 = 0/3, step 2 = 1/4, step 3 = 2 no Q,step 4 = 2& Q.

When the hand being relayed out has not yet shown range (when relays stop in 3♡ (Or below)/when you’ve lost a level. This “doesn’t” apply when you haven’t got much range to show. Mostly with the balanced hands with 3/4♠ .) then the reponses are:

Step 1: Any minimum. Then the lowest step which isn’t game in the set trump suit is RKB again when you give normal responses. Bidding game in the trump suit after that is a sign off. Other bids indicate that the relayer is AWOL :) (Actually other bids should be NNF. Small chance of RKB for a minor then wish to play 4Maj/NT).

Step 2: Maximum hand(13^{+} − 15 or 14^{+} after 1♣ response) 0/3 keycards.

Step 3: Maximum 1/4 keycards.

Step 4: Maximum 2 key cards no Q.

Step 5: Maximum 2 key cards & trump Q.

After using RKB in a suit it is only possible to stop in 3NT/4/5/6 of that suit / 7 of anything/ 6 of a higher suit/6 NT. Other bids are part of the spirals.

The spirals work by ordering the cards you want to ask for in this order: Trump Q, Controls in side suits in length order (A or K), Trump J, Q’s in side suits in length order, then the T of trumps,J’s in side suits (if there is enough room for the last ones).

If the trump Q has been shown then obviously that’s ommited. Also you don’t show controls/Q’s in singleton/void suits. The spiral normally works by counting down the sequence until you find a card you don’t hold and then stopping. If opener skips n steps then responder ignores the first n cards on the list.

These spirals are also used after natural auctions except you order any unknown suits in the order ♣/♢/♡/♠ .

5. Interference

This isn’t normally damaging or indeed very frequent (you’ve already got the opening past one person and you’ve set up your GF so you aren’t hopefully in much trouble). One thing it is a complete memory strain. Just working out which steps you’ve lost.

The idea here is to when possible just ignore it. Following deals with interference in front of the person driving the relays.

After a double (i.e., 1♣−P −2♣−P −2♢− x) then if you don’t want to try passing or redoubling just bid the relay step. Subsequently ignore the double after this. Passing suggests to partner playing the contract redoubled if he’s got “good” length in context. Redouble strongly suggests trying to play the contract redoubled. After either of these then responder can just go back into normal relays if the length is missing :)

After step 1 interference then Double = Relay step, Pass suggests clobbering it to partner :) Then double = penalty , bids as for normal relays (W’out wish to try for penalties).

After step 2 interference then Pass = Relay step then responder can double to show step 1 / bid higher. (i.e., 1♣ − P − 1♡ − 1NT − P − P − x is the same hand as 1♣ − 1♡ − 1♠ − 1NT shows). Double here by responder is a strong suggestion that you try defending. Responder should pull if highly unsuitable showing his hand as naturally as possible.

Over step 3 then as over step 2 but you’ve lost 1 step. This is okay as you can often just drop range showing. Or jam the 2 terminal hand types together(They are normally very similair).

Over higher bids there are no more relays :( Pass is forcing and maybe suggests defending, dble strongly suggest doing that, jumps are splinters; new suits (inc support) are natural.

After interference in front of the person answering relays then bassically you play Pass = step 1, double = step 2 (After pass then relayer doubles again to relay). But we don’t bother trying to “gain” steps.

Over higher bids the bidding is natural.